# First-passage properties of the jump process with a drift. The general case

### v1

@@ 3. Analytic structure of $\phi^{\pm}(\lambda;\rho,s)$ in the $\lambda$-plane
In this section, we treat $\phi^{\pm}(\lambda;\rho,s)$ as functions of $\lambda$ and extend (9) to the complex $\lambda$-plane. We begin by establishing several key properties of the Fourier transform $F(k;\rho)$ given by (11), which serves as a crucial elementary building block. Next, we consider a simple yet illustrative example to show why the analytic continuation of the integrals such as (9) differs from naively applying the definition for $\lambda<0$. Guided by this example, we construct the analytic continuations $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$ of $\phi^{\pm}(\lambda;\rho,s)$ to the $\lambda$-plane. In this section we assume that $s<1$ and $\rho>0$; continuations in the $s$- and $\rho$-planes are addressed in the next sections.
@@ 3.1. Properties of $F(k;\rho)$
The expansion of $F(k;\rho)$ as $k\to 0$ is straightforward. Since all moments of $p(t)$ and $q(M)$ are finite, so are the moments of the effective random walk, and thus the expansion reads:
@@
For the expansion as $k\to\infty$, note that $F(k;\rho)$ is a product of two highly oscillatory integrals, each decaying at large $k$. Consider the second integral, assuming that $p(t)$ is supported on $0,\infty)$. Integration by parts yields
@@
In what follows, we write $\zeta_{-}\equiv\zeta_{-}(\rho)$ omitting the explicit $\rho$-dependence to simplify notation. In cases where the probability distributions decay faster than exponential (e.g., if $p(t)$ or $q(M)$ is half-Gaussian), we set $\zeta_{\pm}=\pm\infty$. Furthermore, when referring to $F(k;\rho)$, we will always assume that $k$ lies within the strip of convergence, i.e., $k=k_{R}+\mathrm{i}k_{I}$ with $k_{R}\in\mathbb{R}$ and $k_{I}\in(\zeta_{-},\zeta_{+})$.
@@
The first simple observation about $F(k;\rho)$ is that within the strip of convergence, we have
@@
Note that (40) implies that there are two branching points of the logarithm at $k=\mathrm{i}\zeta_{1,2}$ (see Fig. 5). In principle, there may be other singularities in the logarithmic term arising from the non-real solutions of (39), but they all lie outside the horizontal strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$. With this we conclude the list of properties of $F(k;\rho)$ that are important for the analysis to come.
Figure 5. The structure of $\log[1-sc(\rho)F(k;\rho)]$ in the complex $k$-plane. The function is analytic in the horizontal strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$ with branching points at $k=\mathrm{i}\zeta_{1,2}$. The edges of the strip $\zeta_{1,2}$ are two real solutions of (39) as shown in Fig. 4. Note that (39) may have other non-real solutions, leading to more singularities in the logarithmic term, but due to (40) these singularities lie in the shaded region.
3.2. A toy example
Before performing an analytic continuation of $\phi^{\pm}(\lambda;\rho,s)$ to the $\lambda$-complex plane, let us consider a simplified version of the problem. Specifically, we define two auxiliary functions
$\psi^{\pm}(\lambda)=\exp\left[-\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}k\,\frac{1}{\lambda\pm\mathrm{i}k}\log\left[(k-\mathrm{i}A)(k+\mathrm{i}B)\right]\right],$ (41)
where $A$ and $B$ are positive real parameters and the integral is understood in terms of the principal value.
The functions $\psi^{\pm}(\lambda)$ resemble $\phi^{\pm}(\lambda;\rho,s)$ in structure, but are sufficiently simple to allow explicit evaluation for all complex $\lambda$. They thus offer an illustrative example in which to expose the difference between a naive extension of (41) and a proper analytic continuation. Moreover, there is a pragmatic motive for this brief detour. Similar integrals will reappear in the analysis of first-passage properties, and it is prudent to compute them here, killing two birds with one well-aimed stone.
The standard method for evaluating the integrals of the form (41) involves extending them into the complex $k$-plane and closing the contour either in the lower or in the upper half-plane. If the contribution from the semicircle vanishes at infinity, the integral is determined by the singularities inside the contour. However, in the form (41) it is not apparent why the contributions from the arcs decay sufficiently fast (faster than $\left|k\right|^{-1}$). For this reason, we modify this representation so that the standard procedure applies.
$\psi^{\pm}(\lambda)=\exp\left[-\frac{1}{4\pi}\int_{-\infty}^{\infty}\mathrm{d}k\left\{\frac{\log\left[(k-\mathrm{i}A)(k+\mathrm{i}B)\right]}{\lambda\pm\mathrm{i}k}+\frac{\log\left[(k+\mathrm{i}A)(k-\mathrm{i}B)\right]}{\lambda\mp\mathrm{i}k}\right\}\right].$ (42)
Rearranging the terms in the integrand we obtain
Applying (45) to the integrals in (44) after algebraic manipulations, we find that
[…]
From (46) it is evident that the analytic continuations $\psi_{\text{a.c.}}^{\pm}(\lambda)$ of the functions $\psi^{\pm}(\lambda)$ from the positive half of the real line $\lambda\in(0;\infty)$ to the full $\lambda$-plane are given by:
Not only do the functions $\psi^{\pm}(\lambda)$, computed in (46) via the original integral expression (41) for arbitrary $\lambda$, differ from their analytic continuations (47), they are in fact discontinuous at $\mathrm{Re}(\lambda)=0$. Having this explicit computation in mind, we now show how to perform the analytic continuation at the level of the integral expression (41).
We first focus on the function $\psi^{+}(\lambda)$. The analytic structure of the integrand for $\psi^{+}(\lambda)$ in the $k$-plane is relatively simple: it has a branch cut from the logarithm and a pole at $k=\mathrm{i}\lambda$. When $\mathrm{Re}(\lambda)\to 0^{+}$, this pole approaches the real axis from above, falling on the integration contour when $\mathrm{Re}(\lambda)=0$ and crossing it to the lower half-plane for $\mathrm{Re}(\lambda)<0$. This is exactly the mechanism responsible for the discontinuity of $\psi^{+}(\lambda)$.
[…]
An important remark here is that the above procedure implicitly assumes that the deformed contour $\mathcal{C}$ does not cross the branch cut of the logarithm. This assumption is however easily lifted. Consider a strip $\mathrm{Im}(k)\in(-a,a)$, where $a=\min(A,B)$. Within this strip, the logarithmic term in the integrand (41) is analytic. If $\left|\,\mathrm{Re}(\lambda)\right|<a$, then the deformed contour lies within $\mathrm{Im}(k)\in(-a,a)$ where there are no singularities. Hence the analytic continuation (50) is valid in this strip. To extend the continuation to the full region $\mathrm{Re}(\lambda)<0$, we combine (50) with (41) to obtain the functional relation
[…]
Clearly (52) and (50) coincide.
@@ 3.3. Continuation in the $\lambda$-plane
The similarity with the example (41) is apparent. There is only one $\lambda$-dependent singularity, i.e., a pole at $k=\pm\mathrm{i}\lambda$. The analytic structure of the logarithm on the other hand depends on the exact form of the function $F(k;\rho)$ and can be rather involved. Fortunately, there is no need to analyze it in detail, and we can focus on the strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$.
@@
Recall that $F(k;\rho)$ is, in general, defined for $\mathrm{Im}(k)\in(\zeta_{-},\zeta_{+})$ hence the lower bound for $\mathrm{Re}(\lambda)$ in (55).
@@
This is the same equation (up to a sign of the argument) as (39). Recall that for the “++” sign, it admits two real solutions $\zeta_{1,2}(\rho,s)$ with $\zeta_{1}<0<\zeta_{2}$ and, due to (40), has no solutions in the strip $\mathrm{Re}(\lambda)\in(\zeta_{1},\zeta_{2})$. Therefore, the rightmost singularity of $\phi^{+}_{\text{a.c.}}(\lambda;\rho,s)$ in the left half-plane is a pole at $\lambda=\zeta_{1}(\rho,s)$. Similarly, the rightmost singularity of $\phi^{-}_{\text{a.c.}}(\lambda;\rho,s)$ is a pole at $\lambda=-\zeta_{2}(\rho,s)$ (see Fig. 7).
@@ 4. Survival probability $S_{\infty}(X_{0})$
Having performed the continuation of $\phi^{\pm}(\lambda;\rho,s)$ into $\lambda$-plane, the next steps are to consider the continuations into $s$- and $\rho$-planes, as the analytic structure in these planes provides us with the asymptotic behaviors of $S_{N}(n\,|\,X_{0})$ and $S_{T}(\tau\,|\,X_{0})$. It is, however, more instructive to start with a more robust quantity: the probability to survive indefinitely $S_{\infty}(X_{0})$. The Laplace transform $\hat{S}_{\infty}(\lambda)$ of $S_{\infty}(X_{0})$ is given by
@@
Finding the exact expression for $\hat{S}_{\infty}(\lambda)$ requires only computing the limit $s\to 1$, $\rho\to 0$ of $\hat{Q}(\lambda;\rho,s)$ rather than performing the full analytic continuation to the $\rho$- and $s$-planes.
In this section we compute the limit in (57) and show that there are indeed two different regimes separated by the critical drift $\alpha_{c}$: the absorption regime, where $S_{\infty}(X_{0})=0$ and the survival regime, where $S_{\infty}(X_{0})>0$. For the survival regime we compute the first two terms of the asymptotic expansion of $S_{\infty}(X_{0})$ when $X_{0}\to 0$ and the rate of exponential decay for $X_{0}\to\infty$.
@@ 4.1. Explicit expression for the Laplace transform
Recall that the analytic structure of integrals entering $\phi_{\text{a.c.}}^{\pm}(\lambda;\rho,s)$ in the $k$-plane described in Sec. 3 (see Fig. 5) is valid for $s<1$ and $\rho>0$. In the limit $s\to 1$ and $\rho\to 0$ the structure changes, as one of the branching points of the logarithm falls onto the real line inducing the singularity at $k=0$. Specifically, if $\langle M-\alpha t\rangle<0$ then $\zeta_{2}=0$, and if $\left\langle M-\alpha t\right\rangle>0$ then $\zeta_{1}=0$ (see Fig. 4). At the same time, as $\lambda\to 0$ the pole $k=-\mathrm{i}\lambda$ in $\phi^{+}_{\text{a.c.}}(\lambda;\rho,s)$ also falls onto the real line at $k=0$. Recall that $\phi_{\text{a.c.}}^{+}(0;\rho,s)$ is one of the factors in the representation (8) for $\hat{Q}(\rho,s\,|\,\lambda)$, hence, in the limit $\rho\to 0$ and $s\to 1$, the coincidence of two singularities at the origin may induce a non-trivial behavior and needs to be dealt with.
Since the problem may arise due to the singularity at $k=0$, we preemptively avoid potential complications by explicitly factoring out the possibly singular behavior from the integral before taking the limit. To simplify the notation we will focus on the case $\lambda>0$. The continuation to $\lambda<0$ can be found straightforwardly from (55). First, we formally rewrite (53) as:
@@
From (67) and (68) the existence of survival and absorption regimes is evident, and so is the critical value of the drift (6) separating these two regimes, i.e., $\alpha_{c}=\langle M\rangle/\langle t\rangle$.
@@ 4.2. Asymptotic behavior
$\hat{S}_{\infty}(\lambda)\underset{\lambda\to\infty}{=}\frac{1}{\lambda}\sqrt{\frac{2\langle M-\alpha t\rangle^{2}}{\left\langle(M-\alpha t)^{2}\right\rangle}}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha t)^{2}\rangle}\right]\right]\\ \times\left(1-\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\;\log\left|1-F(k;0)\right|\right)+O\left(\frac{1}{\lambda^{2}}\right),$ (70)
which, according to Tauberian theorems, translates into
$S_{\infty}(X_{0})\underset{X_{0}\to 0}{\sim}\sqrt{\frac{2\langle M-\alpha t\rangle^{2}}{\langle(M-\alpha t)^{2}\rangle}}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{\frac{1}{2}k^{2}\langle(M-\alpha t)^{2}\rangle}\right]\right]\\ \times\left(1-\frac{X_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\log\left|1-F(k;0)\right|\right).$ (71)
As a consistency check, note that if $\alpha=\alpha_{c}$, then $\langle M-\alpha_{c}t\rangle=0$ and the prefactor in (71) is zero, which agrees with $S_{\infty}(X_{0})=0$ anticipated in (67).
The behavior of $S_{\infty}(X_{0})$ as $X_{0}\to\infty$ can be inferred from the singularity structure of $\hat{S}_{\infty}(\lambda)$ in the $\lambda$-plane. The representation (69) is regular for $\mathrm{Re}(\lambda)>0$ thus, this singularity lies in $\mathrm{Re}(\lambda)\leq 0$. By either keeping track of the analytic structure in (62) and combining the prefactors found in (50) and (55) or, alternatively, by repeating the same arguments as presented in Sec. 3.3 at the level of the representation (68) we find that
@@
As we have argued before in Sec. 3 the denominator has two real solutions at $\lambda=-\zeta_{1,2}$ (see right panel of Fig. 4) and no solutions in between these two values. In the limit $\rho\to 0$ and $s\to 1$ we have $\zeta_{1}=0$, hence in (73) there are two poles: one at $\lambda=0$, and the other at $\lambda=-\zeta_{2}$.
@@
The subleading corrections can be inferred from the singularities of $\hat{S}_{\infty}(\lambda)$. The first of them corresponds to the singularity with the largest real part. In the considered case, this is the pole at $\lambda=-\zeta_{2}(0,1)$. Introducing the notation $R\equiv\zeta_{2}(0,1)$ we write
$S_{\infty}(X_{0})-1\underset{X_{0}\to\infty}{\asymp}e^{-RX_{0}}.$ (76)
@@
As a final remark here, we should mention that the behavior (76) is actually well-known in the financial literature, in which $S_{\infty}(X_{0})$ denotes the probability of the ultimate ruin. In this context, the condition $F(\mathrm{i}R;0)=1$ is essentially equivalent to what is known as the Lundberg fundamental equation [5]. Here we obtain this decay rate as a byproduct, which both serves as a consistency check and demonstrates the power of the representation (8).

### v2

@@ 4. Analytic structure of $\phi^{\pm}(\lambda;\rho,s)$ in the $\lambda$-plane
In this section, we treat $\phi^{\pm}(\lambda;\rho,s)$ as functions of $\lambda$ and extend (9) to the complex $\lambda$-plane. We begin by establishing several key properties of the Fourier transform $F(k;\rho)$ given by (11), which serves as an elementary building block. In particular, we demonstrate that the analytic structure of the integrands defining $\phi^{\pm}(\lambda;\rho,s)$ closely resembles that of the toy example discussed in Section 3: a pole at $k=\mp\mathrm{i}\lambda$ and two branch cuts. The analytic continuations $\phi^{\pm}_{\text{a.c.}}(\lambda;\rho,s)$ to the $\lambda$-plane are thus constructed following the same approach, namely, by tracking the location of the $\lambda$-dependent pole. In this section, we assume that $s<1$ and $\rho>0$; continuations in the $s$- and $\rho$-planes are addressed in the next sections.
@@ 4.1. Properties of $F(k;\rho)$
The expansion of $F(k;\rho)$ as $k\to 0$ is straightforward. Since all moments of $p(t)$ and $q(M)$ are finite, so are the moments of the effective random walk, and thus the expansion reads
@@
For the expansion as $k\to\infty$, we note that $F(k;\rho)$ is a product of two highly oscillatory integrals, each decaying at large $k$. Consider the second integral, assuming that $p(t)$ is supported on $0,\infty)$. Integration by parts yields
@@
In what follows, we write $\zeta_{-}\equiv\zeta_{-}(\rho)$, omitting the explicit $\rho$-dependence to simplify notation. In cases where the probability distributions decay faster than exponential (e.g., if $p(t)$ or $q(M)$ is half-Gaussian), we set $\zeta_{\pm}=\pm\infty$. Furthermore, when referring to $F(k;\rho)$, we always assume that $k$ lies within the strip of convergence, i.e., $k=k_{R}+\mathrm{i}k_{I}$ with $k_{R}\in\mathbb{R}$ and $k_{I}\in(\zeta_{-},\zeta_{+})$.
@@
The first observation about $F(k;\rho)$ is that within the strip of convergence, we have
@@
Note that (56) implies that there are two branch points of the logarithm at $k=\mathrm{i}\zeta_{1,2}$ (see Fig. 7). In principle, there may be other singularities in the logarithmic term arising from the non-real solutions of (55), but they all lie outside the horizontal strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$.
As a final remark, we note that due to (51), the function $F(k;\rho)$ has a saddle point at $k=\mathrm{i}\zeta_{*}$. In its vicinity, we can therefore approximate this function by a quadratic form, which leads to a structure very similar to that of the toy example (29). With this observation, we conclude the list of properties of $F(k;\rho)$ that are important for the analysis to come.
Figure 7. The structure of $\log[1-sc(\rho)F(k;\rho)]$ in the complex $k$-plane. The function is analytic in the horizontal strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$ with branch points at $k=\mathrm{i}\zeta_{1,2}$. The edges of the strip $\zeta_{1,2}$ are two real solutions of (55) as shown in Fig. 6. Note that (55) may have other non-real solutions, leading to more singularities in the logarithmic term, but due to (56) these singularities lie in the shaded region.
@@ 4.2. Continuation in the $\lambda$-plane
The similarity with the example (29) is apparent. There is only one $\lambda$-dependent singularity, i.e., a pole at $k=\pm\mathrm{i}\lambda$. The analytic structure of the logarithm, on the other hand, depends on the exact form of the function $F(k;\rho)$ and can be rather involved. Fortunately, there is no need to analyze it in detail, and we can focus on the strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$.
@@
Recall that $F(k;\rho)$ is, in general, defined for $\mathrm{Im}(k)\in(\zeta_{-},\zeta_{+})$, hence the lower bound for $\mathrm{Re}(\lambda)$ in (59).
@@
This is the same equation (up to the sign of the argument) as (55). Recall that for the “++” sign, it admits two real solutions $\zeta_{1,2}(\rho,s)$ with $\zeta_{1}<0<\zeta_{2}$ and, due to (56), has no solutions in the strip $\mathrm{Re}(\lambda)\in(\zeta_{1},\zeta_{2})$. Therefore, the rightmost singularity of $\phi^{+}_{\text{a.c.}}(\lambda;\rho,s)$ in the left half-plane is a pole at $\lambda=\zeta_{1}(\rho,s)$. Similarly, the rightmost singularity of $\phi^{-}_{\text{a.c.}}(\lambda;\rho,s)$ is a pole at $\lambda=-\zeta_{2}(\rho,s)$ (see Fig. 8).
@@ 5. Survival probability $S_{\infty}(X_{0})$
Having performed the continuation of $\phi^{\pm}(\lambda;\rho,s)$ in the $\lambda$-plane, the next steps are to consider the continuations in the $s$- and $\rho$-planes, as the analytic structure in these planes provides us with the asymptotic behaviors of $S_{N}(n\,|\,X_{0})$ and $S_{T}(\tau\,|\,X_{0})$. It is, however, instructive to start with a more robust quantity: the probability to survive indefinitely $S_{\infty}(X_{0})$. The Laplace transform $\hat{S}_{\infty}(\lambda)$ of $S_{\infty}(X_{0})$ is given by
@@
Finding the exact expression for $\hat{S}_{\infty}(\lambda)$ requires only computing the limit $s\to 1$, $\rho\to 0$ of $\hat{Q}(\lambda;\rho,s)$ rather than performing the full analytic continuation to the $\rho$- and $s$-planes. In the context of the toy example of Section 3, this limit corresponds to the situation in which one of the branch points falls onto the real line.
In this section, we compute the limit in (61) and show that there are indeed two different regimes separated by the critical drift $\alpha_{c}$: the absorption regime, where $S_{\infty}(X_{0})=0$, and the survival regime, where $S_{\infty}(X_{0})>0$. For the survival regime, we compute the first two terms of the asymptotic expansion of $S_{\infty}(X_{0})$ when $X_{0}\to 0$ and the rate of exponential decay for $X_{0}\to\infty$. We then verify the asymptotic expansions numerically, taking both $p(t)$ and $q(M)$ to be inverse-Gaussian distributions.
@@ 5.1. Explicit expression for the Laplace transform
Recall that the analytic structure of integrals entering $\phi_{\text{a.c.}}^{\pm}(\lambda;\rho,s)$ in the $k$-plane described in Section 4 (see Fig. 7) is valid for $s<1$ and $\rho>0$. In the limit $s\to 1$ and $\rho\to 0$ the structure changes, as one of the branch points of the logarithm falls onto the real line inducing the singularity at $k=0$. Specifically, if $\langle M-\alpha t\rangle<0$ then $\zeta_{2}=0$, and if $\left\langle M-\alpha t\right\rangle>0$ then $\zeta_{1}=0$ (see Fig. 6). At the same time, as $\lambda\to 0$ the pole $k=-\mathrm{i}\lambda$ in $\phi^{+}_{\text{a.c.}}(\lambda;\rho,s)$ also falls onto the real line at $k=0$. Recall that $\phi_{\text{a.c.}}^{+}(0;\rho,s)$ is one of the factors in the representation (8) for $\hat{Q}(\rho,s\,|\,\lambda)$, hence, in the limit $\rho\to 0$ and $s\to 1$, the coincidence of two singularities at the origin may induce a non-trivial behavior and needs to be dealt with.
Since the problem may arise due to the singularity at $k=0$, we preemptively avoid potential complications by explicitly factoring out the possibly singular behavior from the integral before taking the limit. To simplify the notation, we focus on the case $\lambda>0$. The continuation to $\lambda<0$ can be found straightforwardly from (59). First, we formally rewrite (57) as:
@@
From (71) and (72) the existence of survival and absorption regimes is evident, and so is the critical value of the drift (6) separating these two regimes, $\alpha_{c}=\langle M\rangle/\langle t\rangle$.
@@ 5.2. Asymptotic behavior
$\hat{S}_{\infty}(\lambda)\underset{\lambda\to\infty}{=}\frac{1}{\lambda}\sqrt{\frac{2\langle M-\alpha t\rangle^{2}}{\left\langle(M-\alpha t)^{2}\right\rangle}}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right]\right]\\ \times\left(1-\frac{1}{\pi\lambda}\int_{0}^{\infty}\mathrm{d}k\;\log\left|1-F(k;0)\right|\right)+O\left(\frac{1}{\lambda^{2}}\right),$ (74)
where $F_{2}(k;\rho)$ is defined in (63). According to Tauberian theorems, (74) translates into
$S_{\infty}(X_{0})\underset{X_{0}\to 0}{\sim}\sqrt{\frac{2\langle M-\alpha t\rangle^{2}}{\langle(M-\alpha t)^{2}\rangle}}\exp\left[-\frac{1}{\pi}\int_{0}^{\infty}\mathrm{d}k\,\mathrm{Im}\left[\frac{1}{k}\log\frac{1-F(k;0)}{1-F_{2}(k;0)}\right]\right]\\ \times\left(1-\frac{X_{0}}{\pi}\int_{0}^{\infty}\mathrm{d}k\;\log\left|1-F(k;0)\right|\right).$ (75)
This is the expansion stated in (17). As a consistency check, note that if $\alpha=\alpha_{c}$, then $\langle M-\alpha_{c}t\rangle=0$ and the prefactor in (75) is zero, which agrees with $S_{\infty}(X_{0})=0$ as in (71).
The behavior of $S_{\infty}(X_{0})$ as $X_{0}\to\infty$ can be inferred from the singularity structure of $\hat{S}_{\infty}(\lambda)$ in the $\lambda$-plane. The representation (73) is regular for $\mathrm{Re}(\lambda)>0$, and thus this singularity lies in $\mathrm{Re}(\lambda)\leq 0$. By either keeping track of the analytic structure in (66) and combining the prefactors found in (39) and (59) or, alternatively, by repeating the same arguments as presented in Section 4.2 at the level of the representation (72) we find that
@@
As we have argued before in Section 4 the denominator has two real solutions at $\lambda=-\zeta_{1,2}$ (see right panel of Fig. 6) and no solutions in between these two values. In the limit $\rho\to 0$ and $s\to 1$ we have $\zeta_{1}=0$, hence in (77) there are two poles: one at $\lambda=0$, and the other at $\lambda=-\zeta_{2}$.
@@
The subleading corrections can be inferred from the singularities of $\hat{S}_{\infty}(\lambda)$. The first of them corresponds to the singularity with the largest real part. Here, this is the pole at $\lambda=-\zeta_{2}(0,1)$. Introducing the notation $R\equiv\zeta_{2}(0,1)$ we write
$1-S_{\infty}(X_{0})\underset{X_{0}\to\infty}{\asymp}e^{-RX_{0}}.$ (80)
@@
As a final remark here, we should mention that the behavior (80) is actually well-known in the financial literature, in which $S_{\infty}(X_{0})$ denotes the probability of the ultimate ruin. In this context, the condition $F(\mathrm{i}R;0)=1$ is essentially equivalent to what is known as the Lundberg fundamental equation [undefd]. Here we obtain this decay rate as a byproduct, which both serves as a consistency check and demonstrates the power of the representation (8).
5.3. Numerical verification
The asymptotic results derived above hold for arbitrary light-tailed distributions $p(t)$ and $q(M)$ with smooth densities. To verify them, it is instructive to consider a specific example. One natural choice would be exponential distributions, which would recover the exact results from [undefal]; this is an exercise we leave to the reader. Since the present paper addresses the case of general distributions, we instead take both $p(t)$ and $q(M)$ to be inverse-Gaussian
$p(t)=p(t;m_{t},\ell_{t})=\sqrt{\frac{\ell_{t}}{2\pi\,t^{3}}}\exp\left[-\ell_{t}\frac{\left(t-m_{t}\right)^{2}}{2m_{t}^{2}t}\right],$ (81)
$q(M)=q(M;m_{M},\ell_{M})=\sqrt{\frac{\ell_{M}}{2\pi\,M^{3}}}\exp\left[-\ell_{M}\frac{\left(M-m_{M}\right)^{2}}{2m_{M}^{2}M}\right].$ (82)
The critical value of the drift is given by $\alpha_{c}=m_{M}/m_{t}$. This choice serves two purposes: first, it clearly falls beyond the Poisson class, providing a genuine test of our general framework; second, while the integrals in (75) still require numerical evaluation, the decay rate in (80) admits a simple closed form.
Figure 9. Comparison of analytical predictions (red dashed lines) with Monte Carlo simulations (black circles) using $10^{6}$ independent trajectory realizations for each value of $X_{0}$. Parameters: $m_{t}=2$, $\ell_{t}=15$, $m_{M}=4$, $\ell_{M}=10$ (so that $\alpha_{c}=2$), and $\alpha=1.8$. Each trajectory includes 10001000 jumps to approximate infinite-time behavior. Left: asymptotic expansion (75) with coefficients computed by numerical integration. Right: exponential decay (80) with rate (85) (logarithmic scale).
For the distributions (81) and (82), expressions (10) and (11) for the effective random walk reduce to
$c(\rho)=\exp\left[\frac{\ell_{M}}{m_{M}}\left(1-\sqrt{1+2\rho\,\frac{m_{M}^{2}}{\alpha\,\ell_{M}}}\right)\right]$ (83)
and
$F(k;\rho)=\frac{1}{c(\rho)}\exp\left[\frac{\ell_{t}}{m_{t}}\left(1-\sqrt{1+2\mathrm{i}\alpha k\frac{m_{t}^{2}}{\ell_{t}}}\right)\right]\\ \times\exp\left[\frac{\ell_{M}}{m_{M}}\left(1-\sqrt{1+2\frac{m_{M}^{2}}{\ell_{M}}\left(\frac{\rho}{\alpha}-\mathrm{i}k\right)}\right)\right].$ (84)
A straightforward calculation then yields the decay rate for the survival probability (80). Specifically, solving the equation $F(\mathrm{i}R;0)=1$ for the smallest positive $R$, we find
$R=2\frac{\ell_{M}\ell_{t}}{m_{M}m_{t}}\left(\frac{\ell_{M}}{m_{M}}+\frac{\ell_{t}}{m_{t}}\right)\frac{\left(m_{M}-\alpha m_{t}\right)}{\left(\ell_{M}+\alpha\ell_{t}\right)^{2}}.$ (85)
Figure 9 compares the analytical predictions with numerical simulations. The agreement confirms both the full asymptotic expansion (75) (left panel) and the exponential decay (80) with rate (85) (right panel).